Suppose the magnitude of the proton charge differs from the magnitude of the electron charge by a mere 1 part in 10^9.

A) What would be the force between two 1.1-mm-diameter copper spheres with 2.0 cm between their centers? Assume that each copper atom has an equal number of electrons and protons. Express your answer to two significant figures and include the appropriate units.

B) Would this amount be detectable?

C) What can you conclude from the fact that no such forces are observed?

The magnitude of the proton charge differs from the magnitude of the electron charge by more than 1 part in 109.
The magnitude of the proton charge equals the magnitude of the electron charge.
The magnitude of the proton charge differs from the magnitude of the electron charge by less than 1 part in 109.
Answer #1

(A) Copper has density ρ = 8.96 g/cm³

and atomic weight 63.546 g/mol

Our spheres each have volume

V = πd³/6 = π * (0.11cm)³ / 6 = 0.000697 cm³

and so have mass

m = ρV = 8.96g/cm³ * 0.000697cm³ = 0.00624 g

which represents n = 0.00624g / 63.546g/mol = 9.83e-5 mol

This means that they each have about

N = 9.83e-5mol * 6.022e23e-/mol = 5.92e19 electrons and protons.

There is an excess charge on each sphere with magnitude

Q = 5.92e19 * 1.602e-19C / 1e9 = 9.48e-9 C

And now for the thrilling conclusion:

F = kQq / d² = 8.99e9N·m²/C² * (9.48e-9C)² / (0.02m)² = 2.0*10^-3 N

(B) No

(C) The magnitude of the proton charge differs from the magnitude of the electron charge by less than 1 part in 109.